OpenStudy (anonymous):

Which shows the expressions in the order they would appear on a number line from least to greatest? answer choices below

OpenStudy (anonymous):

Which shows the expressions in the order they would appear on a number line from least to greatest? square root of eleven then square root of five plus three then eleven divided by seven then two squared then three squared minus the square root of twenty three squared minus square root of twenty then eleven divided by seven then square root of eleven then square root of five plus three, then two squared eleven divided by seven then square root of eleven then two squared then three squared minus square root of twenty, then square root of five plus three square root of five plus three then three squared minus square root of twenty then two squared then eleven divided by seven then square root of eleven

Mehek (mehek14):

so the numbers are √11 √5+3 √ 11/7 2^2 3^- √20?

OpenStudy (anonymous):

yes basically

Mehek (mehek14):

ok √11 = 3.3 (Rounded to the nearest tenth) √5+3 = 2.8 √11/7 = 1.3 2^2 = 4 3^2 - √20 = 4.5